OpenStudy (anonymous):

$89.99 8.5% what would be the total coast

OpenStudy (anonymous):

cost

OpenStudy (anonymous):

i'm trying to figure it out, you should find the actual percentage of 8.5 i don't remember how to convert really. I have an idea but i'm not sure..

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

sure. I'm not sure I can do it though. I see your new to the site though, WELCOME! (:

OpenStudy (anonymous):

:)

OpenStudy (tkhunny):

?? Is that a question? Can yousupply the exact language? Is it: Original Cost is $89.99 and there is 8.5% Tax, so what is the Total Cost?

OpenStudy (anonymous):

@tkhunny yeah i think that's what it was supposed to be

OpenStudy (anonymous):

yes sorry i did not explain 

OpenStudy (tkhunny):

You can calculate the tax separately and add 89.99 * 0.085 = 7.65 89.99 + 7.65 = 97.64 Or, you can calculate the total directly 89.99 * (1 + 0.085) = 89.99 * 1.085 = 97.64